OpenStudy (anonymous):

A package is dropped from a helicopter moving upward at 15 m/s. If it takes 16.0 s before the package strikes the ground, how high above the ground was the package when it was released? (Disregard air resistance.) Note: The answer should be 1000 m

OpenStudy (atlas):

Initial velocity of the package is +15m/s ( I am choosing upwards as +ve direction). Acceleration due to gravity, g is -9.8m/s^2(downwards). You know h =ut +1/2 * at^2 u (initial velocity) = 15 m/s t = 16s a = -9.8 m/s^2 You can get the height h.